### v1

@@ Abstract
A new open-source cluster finding library ”Common Clusterization Library” (CCL) is proposed to describe the clusters production when applied to the transport codes. The new library was applied to the Parton-Hadron-Quantum-Molecular Dynamics approach to be comparable with the established and well known MST and SACA implementations on the exactly same set of the events. The presented procedures were tested with rapidity density distributions and complex observables like $\langle M_{imf}\rangle$ and $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$ at the energies of ALADiN and NA49 experiments. An improvement to the simulated annealing chain is proposed to reduce the computational time. The new stable clusters tracking algorithm for the ”lost” cluster recovering is presented. The results of the ”box-like” coalescence procedure and a mixed method which incorporates the coalescence algorithm steps and the negative binding energy criteria are presented. The difference between the Helium-3 found by the MST-based cluster finding procedure and the coalescence algorithm applied on the exactly same set of events and baryons is shown for the first time.
I Introduction
@@
On order to overcome this limitation a new open-source library ”Common Clusterization Library” (CCL) [ccl_github] is developed on the base of previously presented ”psMST” [Kireyeu:2021igi] library.
@@
The MST-based clusterization will be described in Section II. The Simulated Annealing procedures and their results are covered in Section III. The coalescence approach and ”mixed” procedure are presented in Section V.
@@ III.1 CCL SA description
III.1.1 System Binding Energy
@@
III.1.2 Algorithm Control Parameters
@@ IV Tracking
$E_{\text{bind}}(p)<0\qquad\text{and}\qquad\forall\,c\in C(p):\;t_{\text{coll}}(c)\leq t_{\text{step}},$
where $t_{\rm coll}(c)$ is the current latest collision time recorded for constituent baryon $c$, and $t_{\rm step}$ is the current evaluation time. The $E_{\text{bind}}(p)<0$ condition was chosen for the analysis, but cut in the code can be set to other value or even disabled for some studies. Let $\mathcal{S}=\{s_{1},s_{2},\dots,s_{M}\}$ denote the persistent list of currently registered stable clusters carried over from previous time steps with initial $\mathcal{S}=\varnothing$.
@@
    is the largest previously registered variant of this identity.
    The collision-time advancement condition:
    $\Delta t_{\rm adv}(p,s)\;=\;\exists\,c\in C(p)\cap C(s):\;t_{\text{coll}}^{\text{new}}(c)>t_{\text{coll}}^{\text{old}}(c\mid s).$
    Let
    $\mathit{adv}(p)=\bigvee_{s\in\mathcal{O}(p)}\Delta t_{\rm adv}(p,s)$
    indicate whether any overlapping stored cluster has outdated collision timestamps.
@@
    $A(p)>A(s_{\max})\quad\vee\quad\bigl(A(p)\leq A(s_{\max})\;\wedge\;\mathit{adv}(p)=\text{true}\bigr).$
@@ V.2 Coalescence results
As is seen on the plot, the coalescence approach based on the ”spin-isospin coupling” not only overestimates the experimental data points, but also show the wrong shape of the distribution with local maximum at mid-rapidity $y=0$. The amount of the ”kinetic” deuterons produced in $2\leftrightarrow 3$ reactions through the pion catalyzer is rather small at this energy, however the final result must be the sum of ”kinetic” deuterons and those found by (s)MST or formed via coalescence. The ”sMST + kinetic” sum is closer to the data points and describes the experimental data shape quite well, however, slightly underestimates it. There is no reason to make that similar sum for the probabilistic coalescence as it already overestimate the data.
@@
  • The ”sMST” results also overestimates the data, however, by around 30% at mid-rapidity, and reproduce the experimental shape quite well.
@@ VI Summary
  • The Helium-3 nuclei found by the MST-based cluster finding procedure (”sMST”) and the coalescence algorithm applied on top of the exactly the same set of baryons was presented for the first time and compared to the NA49 experimental data. Despite the overestimation in the measured region, the results favors the ”sMST” procedure by the overall rapidity density shape, what agrees with the previously shown results for deuterons in [Kireyeu:2023spj].
Acknowledgements.
We would like to acknowledge the theoretical support and extensive discussions with Jörg Aichelin and Elena Bratkovskaya, without which this work could not have been completed. We would also like to thank Marcus Bleicher and Jan Steinheimer for the insights on the UrQMD internal coalescence algorithm and its parameters. Also we acknowledge the discussions with Arnaud Le Fèvre on the FRIGA approach. Furthermore, we acknowledge help for the independent cross-checks of the obtained results by Jiaxing Zhao and Vadim Voronyuk. Authors acknowledge JINR for the Young Scientists and Specialists grant number 25-101-08.

### v2

@@ Abstract
A new open-source cluster finding library ”Common Clusterization Library” (CCL) is proposed to describe the clusters production when applied to the transport codes. The new library was applied to the Parton-Hadron-Quantum-Molecular Dynamics approach to be comparable with the established and well known MST and SACA implementations on the exactly same set of the events. The presented procedures were tested with rapidity density distributions and complex observables like $\langle M_{imf}\rangle$ and $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$ at the energies of ALADiN and NA49 experiments. A modification to the simulated annealing chain is proposed to reduce the computational time. The new stable clusters tracking algorithm for the ”lost” cluster recovering is presented. The results of the ”box-like” coalescence procedure and a mixed method which incorporates the coalescence algorithm steps and the negative binding energy criteria are presented. The difference between the Helium-3 found by the MST-based cluster finding procedure and the coalescence algorithm applied on the exactly same set of events and baryons is shown for the first time.
journal: Computer Physics Communications
1 Introduction
@@
On order to overcome this limitation a new open-source library ”Common Clusterization Library” (CCL) [ccl_github] is developed on the base of previously presented ”psMST” [Kireyeu:2021igi] library. The CCL library is distributed under the terms of the ”Let’s Be Scientific” (”LBS”) license which allows for the free code usage and modification, but strictly requires the explicit permission by the authors prior any kind of results publication to avoid the inappropriate use of the code. The full license text is distributed with the library code.
@@
The MST-based clusterization will be described in Section 2. The Simulated Annealing procedures and their results are covered in Section 3. The coalescence approach and ”mixed” procedure are presented in Section 5.
@@ 3.1 CCL SA description
3.1.1 System Binding Energy
@@
3.1.2 Algorithm Control Parameters
@@ 4 Tracking
$E_{\text{bind}}(p)<E^{\text{cut}}_{\text{bind}}\qquad\text{and}\qquad\forall\,c\in C(p):\;t_{\text{coll}}(c)\leq t_{\text{step}},$
where $E^{\text{cut}}_{\text{bind}}\leq 0$ is a configurable threshold (default: any negative value), $t_{\rm coll}(c)$ is the current latest collision time recorded for constituent baryon $c$, and $t_{\rm step}$ is the current evaluation time. Let $\mathcal{S}=\{s_{1},s_{2},\dots,s_{M}\}$ denote the persistent list of currently registered stable clusters carried over from previous time steps with initial $\mathcal{S}=\varnothing$. Prior to processing new candidates, $\mathcal{S}$ is first passed through the ”filtering” procedure: any stored cluster $s\in\mathcal{S}$ is removed if:
  • 1.
    any of its constituents has disappeared, or
  • 2.
    any constituent has undergone a collision after the time when $s$ was registered (i.e., its recorded $t_{\text{coll}}$ is outdated compared to the current global state).
@@
    be the largest previously registered variant of this identity.
@@
    $A(p)>A(s_{\max})$
@@ 5.2 Coalescence results
As is seen on the plot, the coalescence approach based on the ”spin-isospin coupling” not only overestimates the experimental data points, but also show the shape of the distribution with local maximum at mid-rapidity $y=0$, which differs from the experimental data shape. The amount of the ”kinetic” deuterons produced in $2\leftrightarrow 3$ reactions through the pion catalyzer is rather small at this energy, however the final result must be the sum of ”kinetic” deuterons and those found by (s)MST or formed via coalescence. The ”sMST + kinetic” sum is describes the experimental data shape quite well, however, underestimates it. There is no reason to make that similar sum for the probabilistic coalescence as it already overestimate the data.
@@
  • 2.
    The ”sMST” results fits the data perfectly.
@@ 6 Summary
  • 9.
    The Helium-3 nuclei found by the MST-based cluster finding procedure (”sMST”) and the coalescence algorithm applied on top of the exactly the same set of baryons was presented for the first time and compared to the NA49 experimental data. Results favors the ”sMST” procedure.
7 Acknowledgments
We would like to acknowledge the theoretical support and extensive discussions with Jörg Aichelin and Elena Bratkovskaya, without which this work could not have been completed. We would also like to thank Marcus Bleicher and Jan Steinheimer for the insights on the UrQMD internal coalescence algorithm and its parameters.
Also we acknowledge the discussions with Arnaud Le Fèvre on the FRIGA approach.
Furthermore, we acknowledge help for the independent cross-checks of the obtained results by Jiaxing Zhao and Vadim Voronyuk.
Authors acknowledge JINR for the Young Scientists and Specialists grant number 25-101-08.